1. Consider the market supply curve which passes through the intercept and from which the market equilibrium data is known, this is, the price and quantity of equilibrium 𝑷𝑬=πŸ“πŸŽ and 𝑸𝑬=𝟐𝟎𝟎𝟎.

a. Considering those two points, find the equation of the supply.

b. Draw a graph of this line.


1
Expert's answer
2022-03-23T14:02:48-0400

Find the slope

"(0,0) ,(2000,50)"

"\frac{50}{2000}= \frac{1}{40}"

We use the slope to obtain the equation

"\frac{1}{40}= (2000,50),(Q,P)"

"\frac{1}{40}= \frac{P-50}{Q-2000}"

"Q=40P"

The graph of the line is